Find the equation of the tangent and normal to x^2 + 3xy + y^2 = 5 at P(1,1).
Answer by Alan3354(69443)

Find the equation of the tangent and normal to x^2 + 3xy + y^2 = 5 at P(1,1).
---------
Find the 1st derivative
x^2 + 3xy + y^2 = 5
Differentiate implicitly.
2xdx + 3ydx + 3xdy + 2ydy = 0
(2x+3)dx + (2y+3)dy = 0
dy/dx = -(2x+3)/(2y+3) = m, the slope
At (1,1) m = -1
-----
At (1,1), y-1 = -(x-1)
y = -x + 2 is tangent.
--------------
y = x is normal